OpenStudy (anonymous):

A spherical water tank has a diameter of 42 feet. How much water can the tank hold when full?

OpenStudy (anonymous):

10,561π cubic feet 12,348π cubic feet 15,934π cubic feet 21,568π cubic feet

OpenStudy (amonoconnor):

r = d/2 = (42)/2 = 21 V = (4/3)pi(r^3) = pi(4/3)(21^3) = pi(4/3)(9216) = pi(12348) = 12,348pi ft.^3 So... The second choice give. :)

OpenStudy (xcharlizea):

its B :)